Circular No. 7544

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2000 Y1 (TUBBIOLO) R. S. McMillan, Lunar and Planetary Laboratory, reports the discovery

by Andrew F. Tubbiolo of a faint comet with the 0.9-m Spacewatch telescope at Kitt Peak. The object showed a 20′′–30′′ tail on Dec. 16 and 17. The discovery observation is given below; additional observations (including prediscovery observations by LINEAR back to Oct. 24) and parabolic orbital elements (T = 2001 Feb. 6, i = 138^{o}, q = 7.97 AU) are given on MPEC 2000-Y06.

2000 UT  α_{2000}  δ_{2000}  m_{1}
Dec. 16.18043  3^{h}38^{m}09.^{s}56  +20^{o}24′35′′.4  19.3

2000 YA This object, discovered by LONEOS on Dec. 16, will pass ∼ 0.0049

AU from the earth on Dec. 22.28 UT (see MPEC 2000-Y03) at V ∼ 15.

R. P. Binzel and A. S. Rivkin, Massachusetts Institute of Technology; and A. W. Harris, Jet Propulsion Laboratory, report that 0.5–1.0-μm spectra of 2000 YA, obtained on Dec. 17.2 UT using the Palomar 5-m reflector, reveal an S-type asteroid reflectance. For an assumed S-type albedo of 0.2 and H = 23.7, the estimated size is ∼ 50 m.

URSID METEORS 2000 P. Jenniskens, Ames Research Center, supplies a prediction (Jenniskens

and Lyytinen 2000, WGN, submitted) of enhanced activity this year of the meteor stream associated with comet 8P/Tuttle. Notable Ursid outbursts near the time of 8P’s aphelion occurred in 1945 and 1986. Maximum activity is anticipated around Dec. 22.31 UT from material ejected from the comet in 1405. Ejections in 1392 and 1378 could expand this activity over an interval of 4–5 hours.

SUPERNOVA 2000fd IN UGC 291 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum of SN 2000fd (cf. IAUC 7540), obtained by P. Berlind on Dec. 15.14 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova several weeks after maximum light.

2000 December 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green